# Orthogonal roots, quantum Hafnians, and generalized Rothe diagrams

### v1

@@ 4.2. An effective construction
    The rook configuration model described in Example 4.7 (iii) is inspired by a similar model used by Garsia and Remmel in [3] to study a $q$-analogue of the Stirling numbers of the second kind. Garsia and Remmel place rooks on a general Ferrers board, and our model corresponds to the special case where the board is a $a\times b$ rectangle containing $\min(a,b)$ non-attacking rooks, with the residues in our construction corresponding to the positions marked with a circle on the board; see [3, Figure (I.7)].
  • (iii)
    The rook configuration models mentioned in Example 4.7 allow combinatorial proofs of the Poincaré polynomials of the quasiparabolic triples. In particular, it follows from the models that for all the quasiparabolic sets mentioned in the example, the Poincaré polynomials factor into quantum integers (that have combinatorial interpretations). For example, the quasiparabolic sets associated to $T(A_{n},p)$ and $T(A_{n},n+1-p)$ have the same Poincaré polynomial for any $1\leq p\leq\lfloor n+1-p\rfloor$, namely, $\prod_{d=n-2p+2}^{n+1-p}[d]_{q}$.
@@
  1. (i)
    Table 1 gives the complete list of cases for which $T(E_{n},p)$ is a QP-triple, where $n\in\{6,7,8\},1\leq p\leq n$, and $\kappa$ stands for the cardinality of the elements in $\Omega_{U}$ in each case. We have verified that this is the complete list by computation, and we will give a detailed treatment of $T(E_{8},8)$ and $T(E_{7},7)$, two examples that have natural connections to minuscule representations and del Pezzo surfaces, but we do not have a combinatorial model or a conceptual proof of the quasiparabolic structure for all the examples. We also note that by our computation, in the cases where $T(E_{n},p)$ is not a QP-triple, it always happens that $T(E_{n},p)$ is not even a scaled $W_{I}$-set.
    $T(\Gamma,p)$ $|U|$ $|\Omega_{U}|$ $\kappa$ degrees
    $T(E_{6},1)$ 16 40 2 5, 8
    $T(E_{6},2)$ 20 30 4 2, 3, 5
    $T(E_{6},6)$ 16 40 2 5, 8
    $T(E_{7},1)$ 32 120 4 3, 5, 8
    $T(E_{7},2)$ 35 30 7 2, 3, 5
    $T(E_{7},5)$ 35 30 7 2, 3, 5
    $T(E_{7},7)$ 27 45 3 5, 9
    $T(E_{8},1)$ 64 240 8 2, 3, 5, 8
    $T(E_{8},2)$ 64 240 8 2, 3, 5, 8
    $T(E_{8},5)$ 64 240 8 2, 3, 5, 8
    $T(E_{8},6)$ 64 240 8 2, 3, 5, 8
    $T(E_{8},8)$ 56 630 4 5, 9, 14
    Table 1. Table of QP-triples of the form $T(E_{n},p)$
  2. (ii)
    The structures of the QP-triple $T(E_{7},2)$ and the associated quasiparolic set $\Omega(E_{7},2)$ are known by [8, Section 6.2]: the parabolic subgroup $W_{I}$ has type $A_{6}$, the set $\Omega(E_{7},2)$ coincides with the set of positive $7$-roots of $\Phi(E_{7})$ containing no alignments, and $W_{I}\cong S_{7}$ acts transitively on $\Omega(E_{7},2)$ in the standard action. The set $\Omega(E_{7},2)$ is also in a natural bijection with the $30$ inequivalent labellings of the Fano plane. Under the correspondence between the $7$-roots and labelled Fano planes, the generalized Rothe diagram of a labelled Fano plane has a simple combinatorial description: a residue is a triple of non-collinear points such that each point has a higher label than the point collinear with the other two. For example, the unique $W_{I}$-minimal labelling
    $\{136,145,127,235,246,347,567\}$
    is the only labelling with no residues, and the unique $W_{I}$-maximal labelling
    $\{123,145,246,257,347,356,167\}$
    has seven residues: $267$, $357$, $367$, $456$, $457$, $467$, and $567$. The generalized Rothe diagram of each labelling can be interpreted as the set of non-positions in the context of games arising from the Steiner system $S(2,3,7)$; see [11].
  3. (iii)
    For the triple $T(E_{8},2)$, the results of [8, Section 6.3] show that the corresponding subgroup $W_{I}$ has type $A_{7}$, the set $\Omega(E_{8},2)$ coincides with the set of positive $8$-roots of $\Phi(E_{8})$, and $W_{I}\cong S_{8}$ acts transitively on $\Omega(E_{8},2)$ in the standard action. The set $\Omega(E_{8},2)$ is in a natural bijection with the set of $240$ inequivalent labellings of the Fano plane by the set $\{1,2,\ldots,8\}$, as well as with the 240 packings of the projective space $PG(3,2)$. The generalized Rothe diagram of each $8$-root in $\Omega(E_{8},2)$ has a characterization similar to the one in (ii), although the details are more complicated. This particularly intricate example is the subject of the paper [6].
@@ 5. The QP-triples $T(E_{8},8)$ and $T(E_{7},7)$
from Example 4.9 in detail. The sets $U(E_{8},8)$ and $U(E_{7},7)$ have interesting natural connections to minuscule representations and del Pezzo surfaces. We give a non-computational proof that $T(E_{8},8)$ is a QP-triple in Section 5.1, and we use this result to deduce that $T(E_{7},7)$ is a QP-triple in Section 5.2. In Section 5.3, we show that the Pfaffian of $T(E_{7},7)$ coincides with the invariant polynomial associated to a 27-dimensional minuscule representation of type $E_{6}$, and we also describe the generalized Rothe diagrams of the elements of $\Omega(E_{7},7)$, which we expect may have applications to Schubert varieties or del Pezzo surfaces.
@@ 5.1. $T(E_{8},8)$
As we will see in Corollary 5.4, the elements of $\Omega_{U}$ turn out to be precisely the features in $\Phi$ whose elements all come from $U_{1}$, and they can also be characterized as the intersections of the form $Q=R\cap U_{1}$ for the $8$-roots $R$ of $\Phi$ that do not contain $\theta$. Our strategy for proving the quasiparabolic structure of $\Omega_{U_{1}}$ is to leverage the quasiparabolic structure of the set $\Omega_{\Phi_{+}}$ of the $8$-roots, so inclusions of the form $Q\subseteq R$ where $Q\in\Omega_{U_{1}}$ and $R\in\Omega_{\Phi_{+}}$ will play an important role in our arguments, as will chains of the form $Q\subseteq H\subseteq R$ where $R$ is an $8$-root, $H\in D_{6}(R)$, and $Q$ is a feature in $R$. More specifically, it will be useful to study the relationship between the residues of $Q,H,$ and $R$ with various $8$-heights in such a chain. We will do so in lemmas 5.25.3 and 5.55.8.
@@
$\mathrm{Res}^{(i)}(H):=\text{\rm Res}(H)\cap U_{i}\text{\; and \;}\rho^{({i})}(H)=\lvert\mathrm{Res}^{(i)}(H)\rvert$.
@@
    We have $\mathrm{Res}^{(2)}(R)=\{\theta\}$ and $\theta\in\text{\rm Res}_{\Phi_{+}}(H)$.
@@
    If $\lvert H\cap U_{0}\rvert<4$, then $\mathrm{Res}^{(0)}(H)=\emptyset$.
@@
    If $\lvert H\cap U_{1}\rvert<4$, then $\theta\notin\Xi$ and $\mathrm{Res}^{(2)}(H)=\emptyset$.
@@
We prove (ii) by proving its contrapositive. Suppose that $\mathrm{Res}^{(0)}(H)$ contains a root $\gamma$. Applying Proposition 4.1 (i) to $\Xi=\text{\rm Span}(H)\cap\Phi$ shows that $\gamma$ is the highest or second highest root in a $D_{4}$ subsystem $\Psi\subseteq\Xi$ corresponding to some feature $Q\subseteq H$. Each simple root in a $D_{4}$ subsystem appears with nonzero coefficient in each of the two highest roots, so it follows from the assumption $\gamma\in U_{0}$ that $\Psi\subseteq U_{0}$. In particular, the four elements of $Q$ lie in $U_{0}$, so $\lvert H\cap U_{0}\rvert\geq 4$. The conclusion of (ii) now follows.
@@
in $\Xi$, which implies that each of the four roots on the right hand side has $8$-height 11. This contradicts the assumption that $|H\cap U_{1}|<4$, so we have $\theta\notin\Xi$. It follows that $\theta\notin\text{\rm Res}(H)$, so we have $\mathrm{Res}^{(2)}(H)=\emptyset$, proving (iii). ∎
@@
    In the setting of (i), we have $\theta\in\text{\rm Res}(Q_{1})$. We also have
    $\mathrm{Res}^{(2)}(R)=\{\theta\},\quad\mathrm{Res}^{(1)}(R)=\text{\rm Res}_{U_{1}}(Q_{1}),\ \text{and}\quad\mathrm{Res}^{(0)}(R)=\text{\rm Res}(Q_{0}).$
@@
The last paragraph shows that $\theta$ lies in the $D_{4}$-subsystem $\Phi\cap\text{\rm Span}(Q_{1})$. Since $\theta\in\text{\rm Res}_{\Phi_{+}}(Q_{1})$ by Lemma 5.2 (i), it follows that $\theta\in\text{\rm Res}(Q_{1})$. This proves the first assertion of (ii).
The equation $\mathrm{Res}^{(2)}(R)=\{\theta\}$ holds by Lemma 5.2 (ii). The fact that $Q_{1}\subseteq R$ implies that $\mathrm{Res}^{(1)}(R)=\text{\rm Res}_{\Phi_{+}}(R)\cap U_{1}\subseteq\text{\rm Res}_{U_{1}}(Q_{1})$. Conversely, we have $\text{\rm Res}_{U_{1}}(Q_{1})\subseteq\text{\rm Res}_{\Phi_{+}}(R)\cap U_{1}$ because $s_{\beta}(\gamma)>0$ for all $\beta\in Q_{0}\subseteq U_{0}$ and $\gamma\in U_{1}$, so we have $\mathrm{Res}^{(1)}(R)=\text{\rm Res}_{U_{1}}(Q_{1})$. Finally, we have $\text{\rm Res}(R)=\bigcup_{Q\in D_{4}(R)}\text{\rm Res}(Q)$ by Proposition 4.1 (i) and therefore $\mathrm{Res}^{(0)}(R)=\bigcup_{Q\in D_{4}(R)}\mathrm{Res}^{(0)}(Q)$. Every feature $Q$ other than $Q_{0}$ appearing in the union satisfies $\lvert Q\cap U_{0}\rvert<4$ and thus $\mathrm{Res}^{(0)}(Q)=\emptyset$ by Lemma 5.2 (ii), so it follows that $\mathrm{Res}^{(0)}(R)=\mathrm{Res}^{(0)}(Q_{0})$. This completes the proof of (ii). ∎
@@
Since every simple root of $\Xi$ appears in the decomposition of $\psi$ into simple roots and $\psi$ and $\beta_{1}$ have the same $8$-height, it follows that all simple roots of $\Xi$ other than $\beta$ lie in $U_{0}$, proving (iii).
@@
    All residues of $Q$ have $8$-height 1, so that $\text{\rm Res}(Q)=\mathrm{Res}^{(1)}(Q)$.
@@
    $\mathrm{Res}^{(2)}(H)=\{\theta\},\quad\mathrm{Res}^{(1)}(H)=\text{\rm Res}_{U_{1}}(Q),\ \text{and}\quad\mathrm{Res}^{(0)}(H)=\text{\rm Res}(Q_{0}).$

### v2

@@ 4.2. An effective construction
The rook configuration model described in Example 4.7 is inspired by a similar model used by Garsia and Remmel in [5] to study a $q$-analogue of the Stirling numbers of the second kind. Garsia and Remmel place rooks on a general Ferrers board, and our model corresponds to the special case where the board is a $a\times b$ rectangle containing $\min(a,b)$ non-attacking rooks, with the set $\mathscr{D}_{U}(R)$ in our construction corresponding to the positions marked with a circle on the board; see [5, Figure (I.7)].
Example 4.9 (Type $D$).
For all integers $n\geq 4$ and $1\leq p\leq n$, the triple $T(D_{n},p)$ is a QP-triple whenever we have $1\leq p\leq\lfloor n/2\rfloor$, but not when $\lfloor n/2\rfloor<p<n-2$. However, if we relax the setting of Definition 4.6, then the triple
$T^{\prime}(\Gamma,p)=(W(D_{n}),S(D_{n})\setminus\{s_{p}\},U^{\prime}_{p}:=U(D_{n},p)\cap U(D_{n},n)),$
forms a QP-triple for all $1\leq p\leq n-2$. Note that the QP-triple $(W,I,U)$ from Theorem 3.14 is a special case of the triple $T^{\prime}(\Gamma,p)$ from (v): it is $T^{\prime}(D_{2k},k)$.
We note that for all the QP-triples mentioned in the last paragraph, the maximum-cardinality orthogonal sets $R$ and their generalized Rothe diagrams can be interpreted in terms of rook configurations as in Example 4.7, but on a $p\times(n-p)$ board and using the rules of the Rothe diagrams of Theorem 3.14, in which the attacks come from above and to the left. As with Example 4.7, the rook configuration model can be used to give an elementary combinatorial proof of the factorization of the Poincaré polynomial, but this comes with the caveat that a position on the board may now correspond to either one of two roots $\varepsilon_{i}\pm\varepsilon_{j}$. The fact that $T(D_{n},p)$ fails to be a QP-triple when $\lfloor n/2\rfloor<p<n-2$ has to do with this caveat and the fact that in these cases $R$ does not contribute a rook to every row. By contrast, intersecting $U(D_{n},p)$ with $U(D_{n},n)$ to produce $U^{\prime}_{p}$ eliminates this complication for $T^{\prime}(D_{n},p)$: all elements of $U(D_{n},n)$ are of the form $\varepsilon_{i}+\varepsilon_{j}$, so the positions on the board are now in a natural one-to-one correspondence with the elements of $U^{\prime}_{p}$, which is why $T^{\prime}(D_{n},p)$ does form a QP-triple for all $1\leq p\leq n$ in a similar way to Example 4.7.
Example 4.10 (Type $D$).
For any integer $k\geq 2$, the triples $T(D_{2k},2k)$, $T(D_{2k},2k-1)$, $T(D_{2k-1},2k-1)$, and $T(D_{2k-1},2k-2)$ are all QP-triples. Note that they are different from the QP-triples in Example 4.9.
Of the above four triples, the first triple $T(D_{2k},2k)$ is precisely the triple $(W,I,U)$ appearing in Theorem 3.11. The other three triples give rise to essentially the same quasiparabolic set as $T(D_{2k},2k)$, in the sense that their associated quasiparabolic sets $\Omega_{U(\Gamma,p)}$ all admit canonical level-preserving bijections to the quasiparabolic set $\Omega(D_{2k},2k)$. This bijection is induced by the symmetry of the Dynkin diagram for $\Omega(D_{2k},2k-1)$. For the case of $\Omega=\Omega(D_{2k-1},2k-1)$, each element of $\Omega$ corresponds to a (non-perfect) matching of $\{1,2,\dots,2k-1\}$ into (k1)(k-1) pairs and one unmatched point, and the canonical bijection from $\Omega$ to $\Omega(D_{2k},2k)$ sends such an element to the $2k$-root in $\Omega(D_{2k},2k)$ corresponding to the unique perfect matching containing those (k1)(k-1) pairs as blocks.
Example 4.11 (Type $E$).
The last seven rows of Table 2 give the complete list of cases for which $T(E_{n},p)$ is a QP-triple for some $n\in\{6,7,8\}$ and $1\leq p\leq n$. We have verified that this is the complete list by computation, and we will give a detailed treatment of the triples $T(E_{7},7)$ and $T(E_{8},8)$, two examples that have natural connections to minuscule representations and del Pezzo surfaces. We also note that by our computation, in the cases where $T(E_{n},p)$ is not a QP-triple, it always happens that $T(E_{n},p)$ is not even a scaled $W_{I}$-set.
Remark 4.12.
  1. (i)
    The structures of the QP-triple $T(E_{7},2)$ and the associated quasiparabolic set $\Omega(E_{7},2)$ are known by [12, Section 6.2]: the parabolic subgroup $W_{I}$ has type $A_{6}$, the set $\Omega(E_{7},2)$ coincides with the set of positive $7$-roots of $\Phi(E_{7})$ containing no alignments, and $W_{I}\cong S_{7}$ acts transitively on $\Omega(E_{7},2)$ in the standard action. To each root $\alpha\in U(E_{7},2)$ we may naturally associate a triple $t_{\alpha}$ of integers from the set $7]:=\{1,2,...,7\}$, and $\Omega(E_{7},2)$ can be naturally identified with the $30$ inequivalent labellings of the Fano plane using the set $[7]$. It can be shown that a root $\alpha\in U(E_{7},2)$ dominates a labelling $L$ if and only if the elements of $t_{\alpha}$ do not form a line in $L$ and each element in $t_{\alpha}$ is larger than the unique element collinear with the other two in $L$. For example, the unique $W_{I}$-minimal labelling
    $\{136,145,127,235,246,347,567\}$
    is the only labelling dominated by no root, and the unique $W_{I}$-maximal labelling
    $\{123,145,246,257,347,356,167\}$
    is dominated by seven roots: $267$, $357$, $367$, $456$, $457$, $467$, and $567$. The generalized Rothe diagram of each labelling can be interpreted as the set of non-positions in the context of games arising from the Steiner system $S(2,3,7)$; see [16].
  2. (ii)
    For the QP-triple $T(E_{8},2)$, the results of [12, Section 6.3] show that the corresponding subgroup $W_{I}$ has type $A_{7}$, the set $\Omega(E_{8},2)$ coincides with the set of positive $8$-roots of $\Phi(E_{8})$, and $W_{I}\cong S_{8}$ acts transitively on $\Omega(E_{8},2)$ in the standard action. The set $\Omega(E_{8},2)$ is in a natural bijection with the set of $240$ inequivalent labellings of the Fano plane by the set $\{1,2,\ldots,8\}$, as well as with the 240 packings of the projective space $PG(3,2)$. The generalized Rothe diagram of each $8$-root in $\Omega(E_{8},2)$ has a characterization similar to the one in (i), although the details are more complicated. This particularly intricate example is the subject of the paper [8].
4.3. Binary matroids and symmetric spaces
The goal of this subsection is to explain several intriguing connections that the QP-triples from Table 1 and Table 2 have with matroids and with symmetric spaces. We refer the reader to [7, Chapter 1 and Chapter 2] for basic definitions concerning matroids. For background on symmetric spaces, including Cartan’s classification of them, we refer the reader to [2, Chapter 28] and [14, Chapter X].
Let $T=(W,I,U)$ be a QP-triple described in Table 1 or Table 2. It turns out that every $R\in\Omega_{U}$ gives rise to a matroid, $M_{R}$, whose ground set is $R$ and whose set of flats is given by the set
(4.1) $\{R\cap\left(\bigcap_{R^{\prime}\in Y}R^{\prime}\right):Y\subseteq\Omega_{U}\}.$
Furthermore, the isomorphism type of $M_{R}$ depends only on the triple $T$ and not on $R$.
The isomorphism type of the matroid $M_{R}(R\in\Omega_{U})$ is given in the last columns of Tables 1 and 2, whose notation we now explain below. For nonnegative integers $j\leq k$, we use $U_{j,k}$ to denote the uniform matroid with ground set $[k]$ and rank $j$, where a subset of the ground set is independent if and only if it has size at most $j$. The notation $U^{(2)}_{j,k}$ stands for the doubling of $U_{j,k}$, where the ground set and independent sets are given by the doublings of those of $U_{j,k}$. Here, by a doubling of a set $A\subseteq[k]$ we mean the set $A\cup\{a^{\prime}:a\in A\}$, where $1,1^{\prime},2,2^{\prime},...,k,k^{\prime}$ are understood to be $2k$ distinct elements. For each positive integer $n$, the matroid $AG(n,2)$ is the $n$-dimensional affine space over ${\mathbb{F}}_{2}$, where the nonempty flats are given by the affine subspaces. Finally, the matroid $PG(n,2)$ is the $n$-dimensional projective space over ${\mathbb{F}}_{2}$, with the nonempty flats given by the projective subspaces. All of these matroids are examples of binary matroids, in the sense that they are representable over the field ${\mathbb{F}}_{2}$. More details about $U_{j,k},AG(n,2),PG(n,2),$ and binary matroids can be found in sections 1.2, 5.1, 5.4, and 6.3 of [7], respectively.
Remarkably, we can use the matroid $M_{R}$ to predict the degrees of $U$ in all cases, in the following way. If we list the possible sizes for the flats of $M_{R}$ as $k_{0},k_{1},...,k_{r}$, in increasing order, then it turns out that each flat of size $k_{i}$ is a subset of precisely $d_{i}$ elements of $\Omega_{U}$ for some number $d_{i}$ that depends only on $i$ and not on the specific choice of the flat (or $R$), and that the sequence $(d_{0}/d_{1},d_{1}/d_{2},\ldots,d_{r-1}/d_{r})$ produces the sequence of degrees of the QP-set.
@@ 4.3. Binary matroids and symmetric spaces
For the triple $(W,S,\Phi_{+})$ of type $E_{8}$ and for each $8$-root $R\in\Omega_{\Phi_{+}}$, the possible sizes for the flats of $M_{R}$ are $0,1,2,4$ and $8$, and each flat of these sizes is contained in precisely $d_{0}=2025,d_{1}=135,d_{2}=15,d_{3}=3,$ and $d_{4}=1$ element(s) of $\Omega_{U}$, respectively; it is possible to deduce the numbers $d_{1},d_{2},d_{3}$, and $d_{4}$ from Remark 2.1. The quotients $d_{i}/d_{i+1}$ are $2025/135=15$, $135/15=9$, $15/3=5$, and $3/1=3$, which agree with the degrees of $U$ by Remark 4.4 (i).
As is the case in examples 4.74.11, the assertions from the three paragraphs preceding Example 4.13 can be checked using combinatorial arguments in types $A$ or $D$ and via computer verification (or by careful case-by-case analysis) in type $E$. However, we do not have a conceptual explanation for why the matroids can predict the degrees of $U$ via the quotients $d_{i}/d_{i+1}$. Indeed, we found it surprising that the predictions work for all our examples, because the degrees arise from the level generating function of $\Omega_{U}$ but the definition of the matroids $M_{R}$ does not directly involve generalized Rothe diagrams at all. It would be interesting to know whether the intersections in the set given by (4.1) have any inherent connections with generalized Rothe diagrams.
We now restrict our attention to the QP-triples $T(\Gamma,p)$ described in Table 2 and discuss their connection with symmetric spaces. Recall that symmetric spaces are classified by Cartan, who reduced the classification to that of real simple Lie algebras ([3], [14, Chapter X]). To each irreducible symmetric space $G/K$ one may associate a so-called symmetric pair $({\mathfrak{g}},{\mathfrak{k}})$, which is a pair of real semisimple Lie algebras satisfying a Cartan decomposition ${\mathfrak{g}}={\mathfrak{k}}\oplus{\mathfrak{p}}$, where $\mathfrak{k}$} is the fixed-point subalgebra of a Lie algebra involution $\theta$ and $\mathfrak{p}$} is the (1)(-1)-eigenspace of $\theta$. The pair $({\mathfrak{g}},{\mathfrak{k}})$ is not necessarily unique, but the isomorphism types of $\mathfrak{g}$} and $\mathfrak{k}$} are uniquely determined by the symmetric space $G/K$ ([14, Theorem 6.1]). The roots of $\mathfrak{g}$} whose root spaces lie in $\mathfrak{p}$} are the noncompact roots of $({\mathfrak{g}},{\mathfrak{k}})$ ([14, Section VIII.7]).
Using Cartan’s classification of symmetric spaces and symmetric pairs ([14, Table X.V]), it can be checked case by case that for every symmetric space $G/K$ whose associated symmetric pair $({\mathfrak{g}},{\mathfrak{k}})$ consists of two simply-laced Lie algebras of the same rank, we may choose $\mathfrak{k}$} in such a way that there is a QP-triple $(W,I,U)=T(\Gamma,p)$ from Table 2 with the property that $\mathfrak{g}$} is of type $\Gamma$, $U=U(\Gamma,p)$ consists precisely of the noncompact roots of $({\mathfrak{g}},{\mathfrak{k}})$, and we have $I=S\cap W^{\prime}$ where $W^{\prime}$ is the Weyl group of $\mathfrak{k}$}. Moreover, every QP-triple $(W,I,U)=T(\Gamma,p)$ in Table 2 satisfies the above property for a suitable symmetric pair $({\mathfrak{g}},{\mathfrak{k}})$ of equal-rank simply-laced Lie algebras associated to a symmetric space.
Example 4.14.
The QP-triple $(W,I,U):=T(E_{8},8)=(W(E_{8}),S(E_{8})\setminus\{s_{8}\},U(E_{8},8))$ corresponds to the symmetric space of Cartan type EIX. In this case, $\mathfrak{g}$} is the Lie algebra of type $E_{8}$, the positive roots of $\mathfrak{p}$} are the roots in $U$, i.e., the positive roots in which the simple root $\alpha_{8}$ appears with odd coefficient. The positive roots of $\mathfrak{k}$} are the positive roots in which the simple root $\alpha_{8}$ appears with even coefficient, which include the highest root $\theta$ of $\Phi$. The Weyl group $W^{\prime}$ of $\mathfrak{k}$} is of type $E_{7}+A_{1}$; the set $I=\{s_{1},s_{2},\ldots,s_{7}\}=S\cap W^{\prime}$ generates the $E_{7}$ component of $W^{\prime}$, while the reflection $s_{\theta}$, which commutes with all elements of $I$, generates the $A_{1}$ component. The common size $\kappa$ of the elements of $\Omega_{U}$ is $4$, which agrees with the rank of the symmetric space in question.
Table 3 details the correspondences between the Cartan types of the symmetric spaces, the equal-rank simply-laced symmetric pairs, and all QP-triples from Table 2. The second to last column of the table gives the rank of the symmetric space, which turns out to be equal to the size $\kappa$ of the elements in $\Omega_{U}$ in all cases. The last column describes the decomposition of $\mathfrak{k}$}, where we consider the one-dimensional abelian Lie algebra $\mathbb{R}$} to be of simply-laced Cartan type $\mathbb{R}$}”. Note that while we have to distinguish the first two rows in Table 2 because the matroid for the QP-triple $T(A_{2k-1},k)$ differs from those associated with the other QP-triples of type $A$, all the QP-triples of type $A$ correspond to a single type of symmetric space, regardless of the parity of the rank of $\Phi$ or the value of $p$. Similarly, the QP-triples of type $D$ correspond only to two types of symmetric spaces, even though they correspond to four distinct types of matroids in Table 2.
$G/K$ $\Gamma$ $p$ $\text{\rm rank}(G/K)$ $\mathfrak{k}$}
AIII $A_{n}$ $1,2,\ldots,n$ $\min(p,n+1-p)$ $A_{p-1}+A_{n-p}+{\mathbb{R}}$
DI $D_{n}$ $1,2,\ldots,\lfloor n/2\rfloor$ $2p$ $D_{p}+D_{n-p}$
DIII $D_{n}$ n1,nn-1,n $\lfloor n/2\rfloor$ $A_{n-1}+{\mathbb{R}}$
EII $E_{6}$ 22 44 $A_{5}+A_{1}$
EIII $E_{6}$ $1,6$ 22 $D_{5}+{\mathbb{R}}$
EV $E_{7}$ $2,5$ 77 $A_{7}$
EVI $E_{7}$ 11 44 $D_{6}+A_{1}$
EVII $E_{7}$ 77 33 $E_{6}+{\mathbb{R}}$
EVIII $E_{8}$ $1,2,5,6$ 88 $D_{8}$
EIX $E_{8}$ 88 44 $E_{7}+A_{1}$
Table 3. Correspondence between symmetric spaces and QP-triples $T(\Gamma,p)$
@@ 5. The QP-triples $T(E_{8},8)$ and $T(E_{7},7)$
from Section 4.2 in detail. The sets $U(E_{8},8)$ and $U(E_{7},7)$ have interesting natural connections to minuscule representations and del Pezzo surfaces. Recall from Section 4.3 that we can verify by computation that $T(E_{8},8)$ corresponds to the matroid $AG(2,2)$ and the symmetric space of type EIX, and that $T(E_{7},7)$ corresponds to the matroid $PG(1,2)$ and the symmetric space of type EVII. Our goal in this section is to give non-computational proofs of four theorems: we show in Theorem 5.9 that $T(E_{8},8)$ is a QP-triple; we deduce in Theorem 5.13 that $T(E_{7},7)$ is a QP-triple; we explicitly describe in Theorem 5.17 the generalized Rothe diagrams of the elements of $\Omega(E_{7},7)$, which we expect may have applications to Schubert varieties or del Pezzo surfaces; and, we prove in Theorem 5.19 that the generalized Pfaffian of $T(E_{7},7)$ coincides with the invariant polynomial associated to a 27-dimensional minuscule representation of type $E_{6}$.
@@ 5.1. $T(E_{8},8)$
As we will see in Corollary 5.4, the elements of $\Omega_{U}$ turn out to be precisely the features in $\Phi$ whose elements all come from $U_{1}$, and they can also be characterized as the intersections of the form $Q=R\cap U_{1}$ for the $8$-roots $R$ of $\Phi$ that do not contain $\theta$. Our strategy for proving the quasiparabolic structure of $\Omega_{U_{1}}$ is to leverage the quasiparabolic structure of the set $\Omega_{\Phi_{+}}$ of the $8$-roots, so inclusions of the form $Q\subseteq R$ where $Q\in\Omega_{U_{1}}$ and $R\in\Omega_{\Phi_{+}}$ will play an important role in our arguments, as will chains of the form $Q\subseteq H\subseteq R$ where $R$ is an $8$-root, $H\in D_{6}(R)$, and $Q$ is a feature in $R$. More specifically, given such a chain it will be useful to study the relationship between the roots with various $8$-heights that dominate $Q,H,$ or $R$. We will do so in lemmas 5.25.3 and 5.55.8.
@@
$\mathscr{D}^{(i)}(H):=\mathscr{D}(H)\cap U_{i}\text{\; and \;}\rho^{({i})}(H)=\lvert\mathscr{D}^{(i)}(H)\rvert$.
@@
    We have $\mathscr{D}^{(2)}(R)=\{\theta\}$ and $\theta\in\mathscr{D}_{\Phi_{+}}(H)$.
@@
    If $\lvert H\cap U_{0}\rvert<4$, then $\mathscr{D}^{(0)}(H)=\emptyset$.
@@
    If $\lvert H\cap U_{1}\rvert<4$, then $\theta\notin\Xi$ and $\mathscr{D}^{(2)}(H)=\emptyset$.
@@
We prove (ii) by proving its contrapositive. Suppose that $\mathscr{D}^{(0)}(H)$ contains a root $\gamma$. Applying Proposition 4.1 (i) to $\Xi=\text{\rm Span}(H)\cap\Phi$ shows that $\gamma$ is the highest or second highest root in a $D_{4}$ subsystem $\Psi\subseteq\Xi$ corresponding to some feature $Q\subseteq H$. Each simple root in a $D_{4}$ subsystem appears with nonzero coefficient in each of the two highest roots, so it follows from the assumption $\gamma\in U_{0}$ that $\Psi\subseteq U_{0}$. In particular, the four elements of $Q$ lie in $U_{0}$, so $\lvert H\cap U_{0}\rvert\geq 4$. The conclusion of (ii) now follows.
@@
in $\Xi$, which implies that each of the four roots on the right hand side has $8$-height 11. This contradicts the assumption that $|H\cap U_{1}|<4$, so we have $\theta\notin\Xi$. It follows that $\theta\notin\mathscr{D}(H)$, so we have $\mathscr{D}^{(2)}(H)=\emptyset$, proving (iii). ∎
@@
    In the setting of (i), we have $\theta\in\mathscr{D}(Q_{1})$. We also have
    $\mathscr{D}^{(2)}(R)=\{\theta\},\quad\mathscr{D}^{(1)}(R)=\mathscr{D}_{U_{1}}(Q_{1}),\ \text{and}\quad\mathscr{D}^{(0)}(R)=\mathscr{D}(Q_{0}).$
@@
The last paragraph shows that $\theta$ lies in the $D_{4}$-subsystem $\Phi\cap\text{\rm Span}(Q_{1})$. Since $\theta\in\mathscr{D}_{\Phi_{+}}(Q_{1})$ by Lemma 5.2 (i), it follows that $\theta\in\mathscr{D}(Q_{1})$. This proves the first assertion of (ii).
The equation $\mathscr{D}^{(2)}(R)=\{\theta\}$ holds by Lemma 5.2 (ii). The fact that $Q_{1}\subseteq R$ implies that $\mathscr{D}^{(1)}(R)=\mathscr{D}_{\Phi_{+}}(R)\cap U_{1}\subseteq\mathscr{D}_{U_{1}}(Q_{1})$. Conversely, we have $\mathscr{D}_{U_{1}}(Q_{1})\subseteq\mathscr{D}_{\Phi_{+}}(R)\cap U_{1}$ because $s_{\beta}(\gamma)>0$ for all $\beta\in Q_{0}\subseteq U_{0}$ and $\gamma\in U_{1}$, so we have $\mathscr{D}^{(1)}(R)=\mathscr{D}_{U_{1}}(Q_{1})$. Finally, we have $\mathscr{D}(R)=\bigcup_{Q\in D_{4}(R)}\mathscr{D}(Q)$ by Proposition 4.1 (i) and therefore $\mathscr{D}^{(0)}(R)=\bigcup_{Q\in D_{4}(R)}\mathscr{D}^{(0)}(Q)$. Every feature $Q$ other than $Q_{0}$ appearing in the union satisfies $\lvert Q\cap U_{0}\rvert<4$ and thus $\mathscr{D}^{(0)}(Q)=\emptyset$ by Lemma 5.2 (ii), so it follows that $\mathscr{D}^{(0)}(R)=\mathscr{D}^{(0)}(Q_{0})$. This completes the proof of (ii). ∎
@@
Since every simple root of $\Xi$ appears in the decomposition of $\psi$ into simple roots and $\psi$ and $\beta_{1}$ have the same $8$-height, it follows that all simple roots of $\Xi$ other than $\beta_{1}$ lie in $U_{0}$, proving (iii).
@@
    All roots that dominate $Q$ have $8$-height 1, so that $\mathscr{D}(Q)=\mathscr{D}^{(1)}(Q)$.
@@
    $\mathscr{D}^{(2)}(H)=\{\theta\},\quad\mathscr{D}^{(1)}(H)=\mathscr{D}_{U_{1}}(Q),\ \text{and}\quad\mathscr{D}^{(0)}(H)=\mathscr{D}(Q_{0}).$